OpenStudy (anonymous):

Find the length of the missing side. Leave your answer in simplest radical form. Answer A. √ 17 B. 113 m C. √ 113 D. √ 71

OpenStudy (anonymous):



OpenStudy (cwrw238):

use pythagoras x^2 = 7^2 + 8^2 so x = sqrt( 49 + 64)

OpenStudy (anonymous):

c